# FlexRAG: A Flexible and Comprehensive Framework for Retrieval-Augmented Generation

### v1

@@ 1 Introduction
: To reduce the engineering burden on researchers, FlexRAG provides complete bilingual documentation and pre-built retrievers available on the Hugging Face Hub, facilitating the rapid implementation of algorithms. Additionally, FlexRAG provides a comprehensive command-line toolkit that facilitates a wide range of tasks, including data preprocessing, retriever construction, system evaluation, and the development of GUI prototypes, as illustrated in Figure 2.
@@ 2.3.3 Assistants
  • OnelineAssistant: An assistant that retrieves information from local knowledge bases and generates responses based on user queries.
@@ 2.4 Evaluation
FlexRAG supports a variety of evaluation metrics for assessing the performance of RAG systems. These metrics can be broadly categorized into three types: retrieval metrics, generation metrics. To ensure the accuracy and reliability of the evaluation results, we employed the widely adopted pytrec_eval [^13] https://github.com/cvangysel/pytrec_eval, sacreBLEU [^14] https://github.com/mjpost/sacrebleu, and Rouge [^15] https://github.com/pltrdy/rouge for metric computation. Additionally, FlexRAG also supports several LLM-as-a-Judge metrics for evaluating the quality of generated responses.
@@ 4 Resource Overhead Analysis
As shown in Figure 6 Under varying batch sizes, FlexRAG consistently exhibits significantly lower overhead compared to FlashRAG in both average wall-clock time and total CPU time, with performance gaps reaching up to an order of magnitude. In terms of memory consumption, FlexRAG also demonstrates substantially lower average and peak memory usage, outperforming the baseline by several times. These results highlight the tangible performance benefits achieved through the incorporation of a memory-mapping mechanism into the system architecture, as well as the optimization of dense index parameters using the ANN-Benchmark toolkit.

### v2

@@ 1 Introduction
To reduce the engineering burden on researchers, FlexRAG provides complete bilingual documentation and pre-built retrievers available on the Hugging Face Hub, facilitating the rapid implementation of algorithms. Additionally, FlexRAG provides a comprehensive command-line toolkit that facilitates a wide range of tasks, including data preprocessing, retriever construction, system evaluation, and the development of GUI prototypes, as illustrated in Figure 2.
@@ 2.3.3 Assistants
  • OnlineAssistant: An assistant that leverages online API calls to access RAG services provided by commercial companies.
@@ 2.4 Evaluation
FlexRAG supports a variety of evaluation metrics for assessing the performance of RAG systems. These metrics can be broadly categorized into two types: retrieval metrics, and generation metrics. To ensure the accuracy and reliability of the evaluation results, we employed the widely adopted pytrec_eval [^13] https://github.com/cvangysel/pytrec_eval, sacreBLEU [^14] https://github.com/mjpost/sacrebleu, and Rouge [^15] https://github.com/pltrdy/rouge for metric computation. Additionally, FlexRAG also supports several LLM-as-a-Judge metrics for evaluating the quality of generated responses.
@@ 4 Resource Overhead Analysis
As shown in Figure 6, under varying batch sizes, FlexRAG consistently exhibits significantly lower overhead compared to FlashRAG in both average wall-clock time and total CPU time, with performance gaps reaching up to an order of magnitude. In terms of memory consumption, FlexRAG also demonstrates substantially lower average and peak memory usage, outperforming the baseline by several times. These results highlight the tangible performance benefits achieved through the incorporation of a memory-mapping mechanism into the system architecture, as well as the optimization of dense index parameters using the ANN-Benchmark toolkit.